OpenStudy (anonymous):

If m = 7x + 7, m = 5y, and m = 140, find the values of x and y

OpenStudy (anonymous):

m=5y, m=140 so 140=5y therefore y=140/5=28. 140=7x+7 therfore 7x=140-7 so 7x=133 therefore x=133/7 so x=19. this ok?

OpenStudy (anonymous):

Can you use the SSS Postulate or the SAS Postulate to prove ?